OpenStudy (anonymous):

A wheel with radius 3m rolls at 18 rad/s.How fast is a point on the rim of the wheel rising when the point is pi/3 radians above the horizontal(and rising)?

OpenStudy (anonymous):

The two variables here are \(\theta\), the angle of rotation and \(x\), the height of the point.

OpenStudy (anonymous):

The relationship between the two would be \(x=\sin\theta\)

OpenStudy (anonymous):

You are given \(\theta\) and \(d\theta / dt\) and asked to find \(dx/dt\).

OpenStudy (anonymous):

@hturkcom Does this analysis seem correct to you?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Wait, there is on mistake. I should have included the radius.

OpenStudy (anonymous):

asking m/s

OpenStudy (anonymous):

Can you correct my mistake?

OpenStudy (anonymous):

I didn't get it.

OpenStudy (anonymous):

What is the height of the point given theta?

OpenStudy (anonymous):

3m

OpenStudy (anonymous):

So you are saying the point is always at 3m even when the wheel is spinning?

OpenStudy (anonymous):

solution supposed to 27.0 m/s.I cannot find this resolution.

OpenStudy (anonymous):

I'm not asking for the solution. I'm asking for the relationship between \(x\) the height of the point, and \(\theta\) the angle of the point.

OpenStudy (anonymous):

\[\pi/3\]

OpenStudy (anonymous):

60

OpenStudy (anonymous):

Nope. Let me draw a picture.

OpenStudy (anonymous):



OpenStudy (anonymous):

What is the equation which describes the relationship between \(x\) and \(\theta\)?

OpenStudy (anonymous):

Do you know trigonometry?

OpenStudy (anonymous):

0+3cos(pi/3)(18rad/sec)

OpenStudy (anonymous):

=27

OpenStudy (anonymous):

Do you understand where that comes from?

OpenStudy (anonymous):

\[\sin \theta= \frac{ y }{ r } ,y=rsin \theta,\frac{ dy }{dt }=\frac{ dr }{ dt } \sin \theta+rcos \theta \frac{ d \theta }{ dt }=0+3\cos(\frac{ \pi }{ 3 })(18)=27\]

OpenStudy (anonymous):

What do you think do I know it?:)